Fermi National Laboratory
 

Operation Summary for April 22, 2011 - April 25, 2011

The Integrated Luminosity for the period from 4/18/11 to 4/25/11 was 49.01 pb-1 (inverse picobarns). NuMI took 3.36E18 of beam during this same time period.

Friday April 22

The day shift began with Operations monitoring Tevatron (TEV) store 8684 (244.35E30) and the Recycler stash (68.9E10), with the Antiproton Source (Pbar) stacking (24.3E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTFB T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations terminated store 8684 at 10:04 AM due to necessary wet engine flywheel repairs at TEV sector B1. Also, cryo system personnel determined that the E1 wet engine needed bearing replacement work as well. Operations reported at 3:37 PM that the cryo system personnel had finished all their repairs and that the cold compressor pump downs had started. Operators prepared D0 for a collision hall access. Booster and Main Injector personnel conducted studies. A Linac RF station (LRF3) tripped off due to a loose driver filament connection; repairs took just under an hour.

During the evening shift, Operations disabled beam to T-978 at 6 PM. Operations put the TEV into HEP shot setup at 7:02 PM and established store 8688 at 10:22 PM with an average initial luminosity (AIL) of 291.66E30. MI Personnel conducted studies.

Saturday April 23

The midnight shift began with Operations monitoring the store 8688 (152.44E30) and the stash (127.5E10), with Pbar stacking (8.9E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to T-978 at 4:10 AM. A Linac RF station (LRF2) tripped off at 6:31 AM and would not reset. Linac personnel came in and investigated. Operations terminated store 8688 at 7:03 AM. Operators prepared D0 for access. A vacuum burst in the P1 beam transfer line caused about 15 minutes of downtime.

During the day shift, D0 came out of access at 10:04 AM. Operations put the TEV through a dry squeeze.

During the evening shift, Linac personnel reported at 5:06 PM that they had completed their work on LRF2; several components needed to be changed. Operations disabled beam to T-978 at 6 PM. Operations put the TEV into shot setup at 7:08 PM and established store 8692 at 8:19 PM with an AIL of 300.02E30. At 10:30 PM, Linac RF tanks 1 and 2 lost their vacuum. Operators called in Linac personnel and prepared the enclosure for access.

Sunday April 24

The midnight shift began with Operations monitoring store 8692, the stash, and the Pbar stack, and with NuMI and MiniBooNE off due to Linac work.

During the midnight shift, Linac personnel reported they had discovered the vacuum leak on the low level RF feed forward system on tank #2. The completed their repairs at 6:50 AM. Linac personnel started bringing up their RF stations. The H- Source tripped off when the turbo controller power supply failed. Linac personnel investigated. Linac Operator reported that a Klystron RF station (KRF5) would not come on after the access.

During the day shift, Linac personnel completed their repairs at 9:32 AM and Operations established beam to the Accelerator Complex. Operations established beam to T-978 at 9:59 AM. At 1:30 PM, an antiproton transverse emittance blow up occurred in the Accumulator. From that point on, when stacks reached 15 mA or greater, the emittance increased, sometimes causing beam problems. Pbar personnel came in and investigated through the end of shift. Operation reported that Booster and Linac RF systems caused some minor downtimes.

During the evening shift, TEV personnel conducted an end-of-store study from 6:15 PM to 8 PM. Operations terminated store 8692 at 8:24 PM. Operators prepared D0 for access at 8:31 PM. The Tevatron quenched while preparing D0 for access. Tevatron recovery followed. D0 came out of access at 9:58 PM.

Monday April 25

The midnight shift began with the TEV recovering from a quench, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations put the TEV into shot setup and established store 8695 at 1:39 AM with an AIL of 293.14E30. Operations reported that the Accumulator continued to see emittance blow ups. Pbar personnel made some changes and reported that they would continue the investigation on Monday morning. Operations established beam to T-978 at 4:20 AM. Operations reported that a Pbar extraction line quadrupole power supply tripped off and caused the loss of an antiproton transfer. Operations called power supply personnel and they replaced a faulty phase relay.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Pbar personnel to work on their antiproton emittance problems, Booster personnel to conduct studies, and MI personnel to conduct coalescing studies.
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/25/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory